OpenStudy (anonymous):

Object A has a positive charge of 4.3 uC while Object B has a positive charge of 3.6 uC and the force between them is 270N. Calculate the distance between these charges.

OpenStudy (anonymous):

\[F = \frac{ kQq }{ r ^{2} }\] so \[r = \sqrt{\frac{ kQq }{ F }} = \sqrt{\frac{ (9*10^{9})*(4.3*10^{-6})*(3.6*10^{-6}) }{ 270 }} = 0.0227m = 22.7mm\]

OpenStudy (anonymous):

\[= 0.0227m = 22.7mm\] For those where my reply is cut off the right hand side of the screen